OpenStudy (anonymous):

how do i do x \mapsto \frac{1}{\sqrt{3 x+13}}?

OpenStudy (zzr0ck3r):

what is mapsto?

OpenStudy (zzr0ck3r):

\[x\mapsto\frac{1}{\sqrt{3x+13}}\]

OpenStudy (zzr0ck3r):

what is the qestion?

OpenStudy (anonymous):

I'm sorry. I have no idea, I just copy pasted the equation. I have to find the domain of the function {1}{\sqrt{3 x+13}}?

OpenStudy (zzr0ck3r):

ahh so we cant divide by 0, and we cant take the square root of a negative number so we have to make sure both of these dont happen we can do this buy finding out what x's make this a true statement 3x+13>0 do you understand why?

OpenStudy (anonymous):

why?

OpenStudy (zzr0ck3r):

well if we let a = 3x+13 we have \[\frac{1}{\sqrt{a}}\]now it should be easy to see why we need a to be greater than 0, and why a cant be 0 do you why a>0 is a must?

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

ok so we said the thing in the square root, which is in the denominator must be greater than 0 so 3x+13 must be > 0

OpenStudy (anonymous):

okay

OpenStudy (zzr0ck3r):

you will basiccally have 3 problems with domain 1) square root of negative numbers 2) dividing by 0 3) taking log of non positive numbers

OpenStudy (zzr0ck3r):

basically*

OpenStudy (zzr0ck3r):

in this example you had both 1) and 2)

OpenStudy (anonymous):

okay

OpenStudy (zzr0ck3r):

when I find domain, I dont look for things that x CAN be, I look for things x CANT be.

OpenStudy (anonymous):

thanks!

OpenStudy (zzr0ck3r):

np